# The Effect of Massive Trans-Neptunian Objects in the Long-term Evolution and Leaking Rates of Neptune’s 3:2 and 2:1 Mean Motion Resonances

### v2

@@ 2.4 Initial Orbital Distribution of Resonant Populations
Regarding inclination, the L7 model shows a considerably higher median for the 3:2 population compared to the SBDB, with values of 18.90 for the L7 model and 10.78 for the SBDB. This is likely due to the SBDB sample exhibiting strong observational biases that preferentially favor the discovery of objects with low inclinations. Conversely, the difference in medians for the 2:1 population is less pronounced, with values of 10.18 for the L7 population and 11.92 for the SBDB.
In the case of the 2:1 MMR, the spread of the observed sample in the semimajor axis is evident due to our ample selection criteria. We made this decision considering that the form of the resonances is not stringent (as suggested by the L7 model distribution) and does not follow simple analytical approximations, especially at low eccentricities and inclinations (see, e.g. R. Malhotra & Z. Chen, 2023). In any case, to remain only with truly resonant particles, we will perform an initial characterization of the population samples shown in Fig. 1.
3 Simulations: Initial filtering of resonant objects
To filter out highly unstable objects in our sample and remain only with potentially long-term stable resonant objects in both populations (theoretical and observed) within the two resonances (3:2 and 2:1), short-term simulations were conducted with the numerical N-body integrator REBOUND (H. Rein & S. F. Liu, 2012). In all our simulations, we used MERCURIUS (H. Rein et al., 2019) with a tolerance parameter of $10^{-8}$, a hybrid integrator that changes between a Wisdom-Holman symplectic mapping and the high-order integrator IAS15 (H. Rein & D. S. Spiegel, 2015) when solving close encounters with massive bodies. The initial characterizing simulations considered the Sun (with the mass of the terrestrial planets and the Moon added to it), the four giant planets, and the massless particles corresponding to our TNO samples. We used a timestep of 36.5 days and an output cadence of 500 yr, with $2\times 10^{4}$, outputs, for a total integration time of 10 Myr.
To identify potential stable members within the resonances, we analyzed for librating behavior the following resonant angles [^7] We also analyzed other possible first-order arguments as given in C. D. Murray & S. F. Dermott (1999), involving the longitudes of pericenter, $\varpi$, and longitudes of the ascending nodes, $\Omega$, both for Neptune and the particle, corresponding to eccentricity and inclination type resonances, however, we did not find other librating arguments for either resonance.:
@@ 3 Simulations: Initial filtering of resonant objects
corresponding to the main resonant angles for the 3:2 and 2:1 MMRs, respectively. In the above equations, $\lambda_{N}$ corresponds to the mean longitude of Neptune, while $\lambda_{p}$ and $\varpi_{p}$ are the mean longitude and longitude of the perihelion of the particle, respectively. Objects with resonant angles librating with total amplitudes below 340 for the entire 10 Myr integration were classified as resonant. Under this criterion, the JPL’s SBDB sample simulations yielded 500 and 97 resonant objects for the 3:2 and 2:1 populations, respectively. For the L7 model data, the numbers are 3161 and 771, respectively, for the 3:2 and 2:1 MMR. This is in excellent agreement with the characterization performed by M. A. Muñoz-Gutiérrez et al. (2019) on the same L7 data. The resulting orbital elements of these simulations are shown in Figure 2, with the same format used in Fig. 1.
@@
Fig. 2 shows the effect of 10 Myr evolution on the particle distributions, where most of the long-term structure has begun to be imprinted. Dynamical sculpting in the current configuration of the solar system has recently been shown to shape some general properties of the observed distributions of resonant populations. However, the finer structure requires additional mechanisms to be reproduced appropriately (see, for instance, S. Balaji et al., 2023).
In our short-term simulations corresponding to the 3:2 MMR (top-left panel in Fig. 2), both the L7 sample and the SBDB particles follow approximately the same semimajor axis vs eccentricity distribution after 10 Myr, having median $a$ values of 39.40 au and 39.41 au (nearly identical to the nominal value of the 3:2 MMR), and median $e$ values of 0.199 and 0.212 for the L7 and the SBDB samples, respectively. For inclination in the same 3:2 resonance (lower left panel), the L7 sample continues to show an overabundance of high-$i$ TNOs, which results from the same overabundance of high-$i$ objects in the initial distribution; median $i$ values after 10 Myr are found to be 18.58 and 10.43 for the L7 and SBDB samples, respectively.
For the 2:1 MMR (right panels of Fig. 2), the small number of particles derived from observations does not allow us to make a significant comparison; however, the L7 sample seems to trace the expected shape for the resonant Twotino population adequately (see e.g. P. Robutel & J. Laskar, 2001; M. A. Muñoz-Gutiérrez et al., 2021). On the other hand, most of the SBDB particles are found within this expected resonant region after 10 Myr. We found median semimajor axis values of 47.73 au and 47.73 au for the L7 and the SBDB samples (both identical to the nominal value of the 2:1 MMR). Median $e$ (top-right panel) and $i$ (bottom-right panel) values are found to be 0.249 and 0.225, and 10.24 and 9.61, again for the L7 particles and the observed SBDB particles, respectively.
@@
4 Results and Discussion
4.1 Long-term simulations of the characterized resonant populations
To explore the secular dynamics of the MMRs, particularly the leaking rate of particles eroding the resonant populations, we run long-term simulations using the subset of librating objects identified in our initial 10 Myr simulations. We ran these simulations from the end of the 10 Myr short-integrations to 4Gyr mark, using a 180 days and an output cadence of $2\times 10^{5}$ yr, with $1.9950\times 10^{4}$, outputs, for a total integration time of $3.99\times 10^{9}$ yr. We use these longer-spaced intervals to maintain a reasonable level of data output. Lastly, we ran an additional short-integration, identical to the one presented in Section 3, from 4.0 Gyr to 4.010 Gyr to have high-frequency data at the end of the simulation for some of the distribution plots presented across Section 4 (similar to the procedure of K. Zhang & B. J. Gladman, 2022). In all sets of simulations, the planetary energy is conserved at values below $10^{-6}$. The same MERCURIUS integrator from the REBOUND package was used for our long-term simulations with the same accuracy parameter as described in Section 3.
@@ 4.1 Long-term simulations of the characterized resonant populations
Figure 3 follows the same format as Figs. 1 and 2, but the distributions show particles that remain in resonance, i.e., they show consistent libration in the final 10 Myr short-integrations performed at the end of the long-term 4 Gyr simulation, with total amplitudes below 340. For the observed SBDB population (blue dots), our simulations produced 396 and 50 resonant objects inside the 3:2 and 2:1 MMRs, respectively. For the L7 model dataset (gray dots), 1919 and 330 objects remained inside the 3:2 and 2:1 MMR, respectively. The distributions of figure 3 show solid sculpting of the 3:2 population (in agreement with, e.g. S. Balaji et al., 2023), where most of the scattered particles located away from the main bulk of the population have already been cleared out. A similar clearing occurs for the 2:1 populations, as we expected.
@@
Regarding inclination, we observe a consistent decrease in the median values for both the SBDB and L7 populations, as well as within both resonances. For the 3:2 MMR, the decline in the median value is less significant compared to that of the 2:1 MMR, where a 66% reduction is noted for the SBDB population. Specifically, the median value decreases from 9.61 at 10 Myr to 6.31 at 4 Gyr. In contrast, the L7 population shows a more pronounced decline in the 3:2 MMR than in the 2:1 MMR, with values decreasing from 18.58 at 10 Myr to 15.65 at 4 Gyr. These observations clearly indicate that particles with larger inclinations are more susceptible to perturbation and ejection from resonances, which aligns with previous findings (e.g. D. Nesvorný et al., 2000; M. S. Tiscareno & R. Malhotra, 2009).
@@
To better visualize and compare the resonant characteristics of our samples and the effects of long-term evolution under the effect of the giant planets only, in Figure 4, we present the libration amplitude and the center of libration of the resonant angle of our studied populations; these angles were determined from the short-term simulations at the beginning (top panels) and end (bottom panels) of the 4 Gyr integration.
@@
In the right panels of Fig. 4, the number of observed SBDB 2:1 MMR particles only allows for a partial comparison with the L7 population. Twotinos are distinguished by libration centers at $90^{\circ}$ (trailing center), $180^{\circ}$ (symmetric center), and $270^{\circ}$ (leading center). For the 2:1 MMR, again no major changes are observed after 4 Gyr on the asymmetric islands, except for a reduction in the number of particles.
@@
For the SBDB, at the beginning of the simulation, the leading, trailing, and symmetric centers comprise approximately 27%, 35%, and 38% of the total population, respectively. By the end of the simulations, although all populations have decreased, the percentages have changed to 26%, 50%, and 24%. The first thing to note is that the symmetric center has evaporated more rapidly than the other two. When examining the evolution of the asymmetric islands, we find that the leading island appears to be evaporating slightly more rapidly than the trailing one; this difference does not seem to be statistically significant. While the most common explanation for the observed asymmetry between the leading and trailing islands is related to the capture efficiency within the 2:1 MMR during Neptune’s early migration, this simulation suggests that the trailing island is more efficient at keeping objects from drifting into the symmetric island.
@@
On the other hand, the L7 model does not delve deep into the nature of the particles trapped in the 2:1 resonance thus neither does it have some evolution history to shape its populations, nor does it look to reproduce the observed ratios between the leading and trailing islands; as such it starts with the leading, trailing, and symmetric centers containing 36%, 34%, and 30% of the total population, respectively; by the end of our simulations these fractions are 30%, 31%, and 39%. The relevance of the symmetric center rises slightly, moving farther away from the observations. This occurs because at the beginning of the integrations, the leading and trailing centers are populated with many objects with relatively large libration (i.e., objects well bound to the resonance but weakly bound to their libration center, with libration amplitudes close to $180^{\circ}$). By the end of the simulation, fewer such objects remain; those that left the leading and trailing centers have migrated deep into the symmetric island, where they require more than 4 Gyr to escape efficiently from the 2:1 resonance. On the other hand, the ratio of leading to trailing populations is consistent with a constant 1:1 ratio, within errors, throughout the entire simulation; however, it is also consistent with a slightly faster evaporation of the leading island, as was observed in the SBDB sample.
@@
These results reinforce the caveat that caution should be exercised when applying the L7 model to study the evolution of specific aspects of the trans-Neptunian region (as mentioned by the authors of the L7 model, e.g. B. Gladman et al., 2012). Real objects tend to occupy the most stable orbits within a given area, whereas the fine details of L7 orbits are assigned randomly, placing them in relatively more ”average” orbits. Additionally, there are many subdivisions within populations that have not been thoroughly considered, meaning that the population ratios will align with the total available phase space rather than with the most stable available phase space. Furthermore, while the long-term integration of sufficiently detailed models should help synthetic models to converge towards a more comprehensive representation of the Solar System, such convergence may not be straightforward and could go through spurious phases before achieving convergence.
@@
The other noticeable feature lies in the fuzziness of the symmetric islands; while the leading and trailing centers appear to be sharp features, the symmetric center seems to have a halo underneath it; these points correspond to objects that are migrating between the symmetric island and one of the other two islands. The details of such transitions result in large libration amplitudes with intermediate libration centers.
@@
Regarding the libration center of the 2:1 population (upper right panel), there is a clear difference for the SBDB population after 4 Gyr; this is due to the substantial evolution of the trailing island, which increases in fractional population by 15%, while the symmetric island has the opposite effect. At the same time, the fractional population of the leading islands remains unchanged. On the other hand, since the biggest change in the fractional populations of the L7 model occurs in the symmetric island, it does not show as a significant jump in the evolution of the L7 model (gray) curves.
Regarding the libration amplitude of the 2:1 MMR (lower right panel), we can see that most of the SBDB population in the asymmetric islands remains protected from ejections from the resonance, going from representing 60% of the initial population to nearly 80% by the end of the integrations. It is interesting to note that the trend of the L7 model moves in the opposite direction of the SBDB, with the population on the asymmetric island decreasing from 70% at the start of the integration to approximately 60% after 4 Gyr. Additionally, there is a notable similarity between the final L7 distribution of libration amplitude and the initial SBDB sample. As seen in Fig. 5, here we can again observe how there are very few objects with libration amplitude in the $100^{\circ}$ to $240^{\circ}$ range; this is because leading orbits limit themselves to $190^{\circ}<\phi<290^{\circ}$ and trailing orbits limit themselves to $70^{\circ}<\phi<170^{\circ}$, while symmetric orbits will always go beyond these limits having a resonant angle that, as a minimum, covers the $60^{\circ}$ to $300^{\circ}$ range.
@@
4.2 The evolution of the most massive Plutinos and Twotinos

### v3

@@ II.4 Initial Orbital Distribution of Resonant Populations
Regarding inclination, the L7 model shows a considerably higher median for the 3:2 population compared to the SBDB, with values of 18.90 for the L7 model and 10.78 for the SBDB. This is likely due to the SBDB sample exhibiting strong observational biases that preferentially favor the discovery of objects with low inclinations. Conversely, the difference in medians for the 2:1 population is less pronounced, with values of 10.18 for the L7 population and 11.92 for the SBDB; this effect is due to an intrinsically tighter inclination distribution of the 2:1 MMR as well as an increased instability for higher inclination objects.
In the case of the 2:1 MMR, the spread of the observed sample in the semimajor axis is evident due to our ample selection criteria. We made this decision considering that the form of the resonances is not stringent (as suggested by the L7 model distribution) and does not follow simple analytical approximations, especially at low eccentricities and inclinations (see, e.g. Malhotra and Chen, 2023). In any case, to remain only with truly resonant particles, we will perform an initial characterization of the population samples shown in Fig. 1.
III Simulations: Initial filtering of resonant objects
To filter out highly unstable objects in our sample and remain only with potentially long-term stable resonant objects in both populations (theoretical and observed) within the two resonances (3:2 and 2:1), short-term simulations were conducted with the numerical N-body integrator REBOUND (Rein and Liu, 2012). In all our simulations, we used MERCURIUS (Rein et al., 2019) with a tolerance parameter of $10^{-9}$, a hybrid integrator that changes between a Wisdom-Holman symplectic mapping and the high-order integrator IAS15 (Rein and Spiegel, 2015) when solving close encounters with massive bodies. The initial characterizing simulations considered the Sun (with the mass of the terrestrial planets and the Moon added to it), the four giant planets, and the massless particles corresponding to our TNO samples. We used a timestep of 36.5 days and an output cadence of 500 yr, with $2\times 10^{4}$, outputs, for a total integration time of 10 Myr.
To identify potential stable members within the resonances, we analyzed for librating behavior the following resonant angles [^7] We also analyzed other possible first-order arguments as given in Murray and Dermott (1999), involving the longitudes of pericenter, $\varpi$, and longitudes of the ascending nodes, $\Omega$, both for Neptune and the particle, corresponding to eccentricity and inclination type resonances, however, we did not find other librating arguments for either resonance.:
@@ III Simulations: Initial filtering of resonant objects
corresponding to the main resonant angles for the 3:2 and 2:1 MMRs, respectively. In the above equations, $\lambda_{N}$ corresponds to the mean longitude of Neptune, while $\lambda_{p}$ and $\varpi_{p}$ are the mean longitude and longitude of the perihelion of the particle, respectively. Objects with resonant angles librating with total amplitudes below 340 for the entire 10 Myr integration were classified as resonant. Under this criterion, the JPL’s SBDB sample simulations yielded 500 and 97 resonant objects for the 3:2 and 2:1 populations, respectively. For the L7 model data, the numbers are 3161 and 771, respectively, for the 3:2 and 2:1 MMR. This is in excellent agreement with the characterization performed by Muñoz-Gutiérrez et al. (2019) on the same L7 data. The resulting orbital elements of these simulations are shown in Figure 2, with the same format used in Fig. 1.
@@
Fig. 2 shows the effect of 10 Myr evolution on the particle distributions, where most of the long-term structure has begun to be imprinted. Dynamical sculpting in the current configuration of the solar system has recently been shown to shape some general properties of the observed distributions of resonant populations. However, the finer structure requires additional mechanisms to be reproduced appropriately (see, for instance, Balaji et al., 2023).
In our short-term simulations corresponding to the 3:2 MMR (top-left panel in Fig. 2), both the L7 sample and the SBDB particles follow approximately the same semimajor axis vs eccentricity distribution after 10 Myr, having median $a$ values of 39.40 au and 39.41 au (nearly identical to the nominal value of the 3:2 MMR, as expected), and median $e$ values of 0.199 and 0.212 for the L7 and the SBDB samples, respectively. For inclination in the same 3:2 resonance (lower left panel), the L7 sample continues to show an overabundance of high-$i$ TNOs, which results from the same overabundance of high-$i$ objects in the initial distribution; median $i$ values after 10 Myr are found to be 18.58 and 10.43 for the L7 and SBDB samples, respectively.
For the 2:1 MMR (right panels of Fig. 2), the small number of particles derived from observations does not allow us to make a significant comparison; however, the L7 sample seems to trace the expected shape for the resonant Twotino population adequately (see e.g. Robutel and Laskar, 2001; Muñoz-Gutiérrez et al., 2021). On the other hand, most of the SBDB particles are found within this expected resonant region after 10 Myr. We found median semimajor axis values of 47.73 au and 47.73 au for the L7 and the SBDB samples (both identical to the nominal value of the 2:1 MMR). Median $e$ (top-right panel) and $i$ (bottom-right panel) values are found to be 0.245 and 0.253, and 10.24 and 9.61, again for the L7 particles and the observed SBDB particles, respectively.
@@
IV Results and Discussion
IV.1 Long-term simulations of the characterized resonant populations
To explore the secular dynamics of the MMRs, particularly the leaking rate of particles eroding the resonant populations, we run long-term simulations using the subset of librating objects identified in our initial 10 Myr simulations. We ran these simulations from the end of the 10 Myr short-integrations to 4Gyr mark, using a 180 days and an output cadence of $2\times 10^{5}$ yr, with 19,950, outputs, for a total integration time of $3.99\times 10^{9}$ yr. We use these longer-spaced intervals to maintain a reasonable level of data output. Lastly, we ran an additional short-integration, identical to the one presented in Section III, from 4.0 Gyr to 4.010 Gyr to have high-frequency data at the end of the simulation for some of the distribution plots presented across Section IV (similar to the procedure of Zhang and Gladman, 2022). In all sets of simulations, the planetary energy is conserved at values below $10^{-6}$. The same MERCURIUS integrator from the REBOUND package was used for our long-term simulations with the same accuracy parameter as described in Section III.
@@ IV.1 Long-term simulations of the characterized resonant populations
Figure 3 follows the same format as Figs. 1 and 2, but the distributions show particles that remain in resonance, i.e., they show consistent libration in the final 10 Myr short-integrations performed at the end of the long-term 4 Gyr simulation, with total amplitudes below 340. For the observed SBDB population (blue dots), our simulations produced 396 and 50 resonant objects inside the 3:2 and 2:1 MMRs, respectively. For the L7 model dataset (gray dots), 1919 and 330 objects remained inside the 3:2 and 2:1 MMR, respectively. The distributions of figure 3 show solid sculpting of the 3:2 population (in agreement with, e.g. Balaji et al., 2023), where most of the scattered particles located away from the main bulk of the population have already been cleared out. A similar clearing occurs for the 2:1 populations, as we expected.
@@
Regarding inclination, we observe a consistent decrease in the median values for both the SBDB and L7 populations, as well as within both resonances. For the 3:2 MMR, the decline in the median value is less significant compared to that of the 2:1 MMR, where the inclination of the SBDB population is reduced to 66% of its initial value. Specifically, the median value decreases from 9.61 at 10 Myr to 6.31 at 4 Gyr. In contrast, the L7 population shows a more pronounced decline in the 3:2 MMR than in the 2:1 MMR, with values decreasing from 18.58 at 10 Myr to 15.65 at 4 Gyr. These observations clearly indicate that particles with larger inclinations are more susceptible to perturbation and ejection from resonances, which aligns with previous findings (e.g. Nesvorný et al., 2000; Tiscareno and Malhotra, 2009).
@@
To better visualize and compare the resonant characteristics of our samples and the effects of long-term evolution under the effect of the giant planets only, in Figure 4, we present the libration amplitude and the center of libration of the resonant angle of our studied populations; these angles were determined from the 10 Myr high-resolution simulations at the beginning (top panels) and end (bottom panels) of the 4 Gyr integrations, where libration centers are determined as the median values, and libration amplitudes are calculated using the maximum and minimum values of their oscillation.
@@
In the right panels of Fig. 4, the number of observed SBDB 2:1 MMR particles only allows for a partial comparison with the L7 population. Twotinos are distinguished by libration centers at $90^{\circ}$ (leading center), $180^{\circ}$ (symmetric center), and $270^{\circ}$ (trailing center). For the 2:1 MMR, again, no major changes are observed after 4 Gyr on the asymmetric islands, except for a reduction in the number of particles.
@@
For the SBDB, at the beginning of the simulation, the leading, trailing, and symmetric centers comprise approximately 35%, 27%, and 38% (34, 26, and 37 particles) of the total population, respectively. By the end of the simulations, although all populations have decreased, the percentages have changed to 50%, 26%, and 24% (25, 13, and 12 particles). The first thing to note is that the symmetric center has evaporated more rapidly than the other two. When examining the evolution of the asymmetric islands, we find that the trailing island appears to be evaporating slightly more rapidly than the leading one; however large, this difference is not statistically significant due to the small number of particles in our 2:1 population. While the most common explanation for the observed asymmetry between the leading and trailing islands is related to the capture efficiency within the 2:1 MMR during Neptune’s early migration (see e.g., Chiang and Jordan, 2002; Chen et al., 2019), this simulation suggests that the trailing island is more efficient at keeping objects from drifting into the symmetric island.
@@
On the other hand, the L7 model does not delve deep into the nature of the particles trapped in the 2:1 resonance thus neither does it have some evolution history to shape its populations, nor does it look to reproduce the observed ratios between the leading and trailing islands; as such it starts with the leading, trailing, and symmetric centers containing 34%, 36%, and 30% (262, 278, and 231 particles) of the total population, respectively; by the end of our simulations these fractions are 31%, 30%, and 39% (104, 102, and 129 particles). The relevance of the symmetric center rises slightly, moving farther away from the observations. This occurs because at the beginning of the integrations, the leading and trailing centers are populated with many objects with relatively large libration (i.e., objects well bound to the resonance but weakly bound to their libration center, with relatively large libration amplitudes close to $140^{\circ}$). By the end of the simulation, fewer such objects remain; those that left the leading and trailing centers have migrated deep into the symmetric island, where they require more than 4 Gyr to escape efficiently from the 2:1 resonance. On the other hand, the ratio of leading to trailing populations is consistent with a constant 1:1 ratio, within errors, throughout the entire simulation; however, it is also consistent with a slightly faster evaporation of the trailing island, as was observed in the SBDB sample.
@@
These results reinforce the caveat that caution should be exercised when applying the L7 model to study the evolution of specific aspects of the trans-Neptunian region (as mentioned by the authors of the L7 model, e.g. Gladman et al., 2012). Real objects tend to occupy the most stable orbits within a given area, whereas the fine details of L7 orbits are assigned randomly, placing them in relatively more ”average” orbits. Additionally, there are many subdivisions within populations that have not been thoroughly considered, meaning that the population ratios will align with the total available phase space rather than with the most stable available phase space. Furthermore, while the long-term integration of sufficiently detailed models should help synthetic models to converge towards a more comprehensive representation of the Solar System, such convergence may not be straightforward and could go through spurious phases before achieving convergence.
@@
The other noticeable feature lies in the fuzziness of the symmetric islands; while the leading and trailing centers appear to be sharp features, the symmetric center seems to have a halo underneath it; these points correspond to objects that are migrating between the symmetric island and one of the other two islands (see also Chen et al., 2019). The details of such transitions result in large libration amplitudes with intermediate libration centers.
@@
Regarding the libration center of the 2:1 population (upper right panel), a clear difference is observed for the SBDB population after 4 Gyr. This is due to the substantial evolution of the leading island, which increases its fractional population by 15%, while the symmetric island has the opposite effect. At the same time, the fractional population of the trailing island remains unchanged. On the other hand, since the biggest change in the fractional populations of the L7 model occurs in the symmetric island, it does not show as a significant jump in the evolution of the L7 model (gray) curves.
Regarding the libration amplitude of the 2:1 MMR (lower right panel), we can see that most of the SBDB population in the asymmetric islands remains protected from ejections from the resonance, going from representing 60% of the initial population to nearly 80% by the end of the integrations. It is interesting to note that the trend of the L7 model moves in the opposite direction of the SBDB, with the population on the asymmetric island decreasing from 70% at the start of the integration to approximately 60% after 4 Gyr. Additionally, there is a notable similarity between the final L7 distribution of libration amplitude and the initial SBDB sample. As seen in Fig. 5, here we can again observe how there is a clear gap in libration amplitude between the symmetric and asymmetric islands; this is because the leading orbits have to limit themselves to instantaneous values of $\phi_{2:1}<180^{\circ}$ (most remain within $30^{\circ}<\phi_{2:1}<170^{\circ}$), trailing orbits have to limit themselves to $\phi_{2:1}>180^{\circ}$ (most remain within $190^{\circ}<\phi_{2:1}<330^{\circ}$); meanwhile, the symmetric island must encompass both other islands (whose exact location depends slightly on $e$ and $i$), resulting in a minimum libration in the $55^{\circ}<\phi_{2:1}<305^{\circ}$ range. The few objects in this gap correspond to objects that are switching between islands.
@@
IV.2 The evolution of the most massive Plutinos and Twotinos